# https://www.vrvis.at

### v1

@@ Abstract
In this work we address flexibility in deep learning by means of transductive reasoning. For adaptation to new tasks or new data, existing methods typically involve tuning of learnable parameters or even complete re-training from scratch, rendering such approaches unflexible in practice. We argue that the notion of separating computation from memory by the means of transduction can act as a stepping stone for solving these issues. We therefore propose PARMESAN (parameter-free memory search and transduction), a scalable transduction method which leverages a memory module for solving dense prediction tasks. At inference, hidden representations in memory are being searched to find corresponding examples. In contrast to other methods, PARMESAN learns without the requirement for any continuous training or fine-tuning of learnable parameters simply by modifying the memory content. Our method is compatible with commonly used neural architectures and canonically transfers to 1D, 2D, and 3D grid-based data. We demonstrate the capabilities of our approach at complex tasks such as continual and few-shot learning. PARMESAN learns up to 370 times faster than common baselines while being on par in terms of predictive performance, knowledge retention, and data-efficiency.
@@
Transduction Memory Correspondence Matching Fast Learning Continual Learning Few-Shot Learning
@@ 1 Introduction
Figure 1: Overview of our parameter-free transduction method for dense prediction tasks. First, a feature extractor $F$ is applied to a query input $x_{q}$ to obtain hidden representations $(h)_{q}^{l}$ for each level $l$. A memory $M$ stores labeled samples along with their hidden representations $(h)_{j}^{l}$. Our approach then performs a hierarchical search in $M$. The search objective is to find both globally and locally similar nodes compared to the query. In each level $l$, we keep the top-$k$ most similar nodes and retrieve their respective child nodes in level $l-1$. Labels $z_{k}$ from most similar memory nodes are retrieved to obtain an initial, raw prediction. Finally, message passing (MP) is applied to obtain a final prediction $y_{q}$. (Simpson images credit: [107, 90, 91, 108])
@@
For training supervised, task-specific deep learning (DL) models, both data and task are usually well-defined in advance. While this setup holds true for many practical problems, a major issue emerges when already existing models require the ability to adapt to both new data and new tasks. Satisfying such a flexibility requirement is non-trivial in practice, adding yet another level of difficulty to a given problem. Although research in fields like continual learning (CL) and few-shot learning (FSL) resulted in many promising approaches, they still suffer from certain restrictions [8, 102] or focus on a specific problem-niche [66]. Moreover, many proposed solutions are often overly complex and require domain experts for deployment and maintenance.
@@
For example, a well established approach is replay-based training in CL, which uses a memory module for knowledge retention [82, 71, 51, 76]. However, replay training turns out to be quite tricky to do in practice due to several reasons. First, models require a high plasticity in order to learn and unlearn continually, which in turn limits their ability to retain knowledge. Forcing the model to be robust, i.e., retain knowledge via replay training, is thus somewhat contradicting the plasticity requirement to learn new concepts. Second, all relevant knowledge has to be incorporated into the model via training, thereby rendering the available memory capacity irrelevant for inference. Third, the balancing of memory samples and new samples for continual training is non-trivial in practice and often leads to the problem of memory overfitting [51, 6, 99]. Finally, replay does not allow to unlearn specific examples easily, e.g., in case they are not needed anymore. Other CL approaches require repeated training from scratch [66], often rendering them infeasible for practical applications.
We argue that simplicity and flexibility are not necessarily mutually exclusive properties of modern DL approaches. In contrast to common, induction-based approaches, where predictions are inferred from general, learned principles, we embrace transduction [21], which is characterized by reasoning from specific training cases to specific test cases.
@@
Our Work: We propose PARMESAN (parameter-free memory search and transduction), an approach that separates computation from memory and extends transduction to complex, high-dimensional labels for dense prediction tasks while at the same time retaining high flexibility. Our method combines several concepts to perform hierarchical correspondence matching with a memory that contains task-specific, labeled data. Rather than requiring to incorporate all knowledge into learnable parameters, the memory in our method extends the capacity of existing models. We further propose a message passing approach to take advantage of correlations within a query sample.
@@
In general, PARMESAN is parameter-free in the sense that it does not have learnable parameters. Instead, learning can be easily and intuitively performed by non-ML experts via memory consolidation, i.e., adding, removing, or modifying memory samples. Our method allows for fast learning and unlearning of individual examples or parts thereof. Unlike many DL approaches for CL and FSL, our method does not require any continual training or fine-tuning. It therefore neither suffers from parameter-induced forgetting and memory overfitting, nor does it require complicated and energy-hungry training by design.
@@
While we apply our method to 2D images in this work, it canonically transfers to other data dimensionalities such as 3D volumes and 1D sequences, as well as being compatible with spatio-temporal data. We do not impose major restrictions on memory size and allow for flexible memory management. Our method can be applied in combination with common DL architectures and is task-agnostic, i.e., transfers to different dense prediction tasks as well as multi-tasking. Our method can be easily integrated into existing DL architectures, thereby enhancing both flexibility and learning speed.
@@
  • We propose a flexible, parameter-free transduction method for dense prediction tasks using hierarchical correspondence matching with a memory module.
  • We further propose a parameter-free message passing approach to take advantage of intra-query correlations.
  • We demonstrate that PARMESAN can be used in combination with common feature extractors.
  • We show that our method can be successfully applied to CL and FSL, being on par with commonly used baselines in terms of predictive performance. At the same time, our method learns up to 370 times faster, has stable knowledge retention properties, and is data-efficient. We demonstrate the capabilities of our approach for semantic segmentation and monocular depth estimation on the Cityscapes and JSRT Chest X-ray datasets.
@@ 2 Related work
Our method combines concepts from the following fields:
@@
Memory Networks: Recently, the field of memory networks [31, 106, 25, 26, 68] experiences renewed interest. When separating computation and memory [26], some parts of an ML model are responsible for computing, others for storing knowledge. With this notion it is possible to add new information easily, while at the same time being able to treat memory contents as variables. Memory consolidation [23, 3] concerns the strategy for deciding which knowledge to keep in the presence of limited memory resources. Sophisticated memory consolidation methods may include, e.g., novelty search [44], where more informative examples replace less informative examples in memory.
@@
Transduction: Recently, transduction methods experience a revival due to their effectiveness on various classification tasks [2, 37]. One of the most prominent transduction methods is based on k-nearest neighbors (k-NN) [16], which aims to find similar examples w.r.t. common features. Transduction via k-NN was demonstrated to work well for image classification [59], where labeled training samples and their hidden representations are stored in memory. Given a query sample, labels of similar samples are retrieved from memory, followed by a majority voting for prediction. Moreover, k-NN in combination with vision transformers [98, 11] yielded promising results for dense out-of-distribution (OOD) detection [20]. However, making k-NN scalable to dense prediction tasks at high resolutions remains underexplored.
@@
Correspondence Matching: Dense correspondence matching [75, 105, 48] [53, 36, 32, 86, 13] is a fundamental problem in computer vision and often acts as a starting point for various downstream tasks such as optical flow, depth-estimation, etc. In this context, promising methods for improving feature correlation have been proposed [94, 34]. Also, methods for exploiting pyramid features have been introduced for object detection [24, 47] and semantic feature matching [95, 111]. Recently, large-scale foundation models [4] have been applied for one-shot semantic segmentation [49]. Many existing matching approaches use a single reference sample with known correspondences, but typically do not scale well to more than one reference sample. Moreover, most methods rely on specific assumptions commonly found in setups with videos or stereo or multi-view images. In contrast, our proposed method allows matching of arbitrary samples and does not require strong assumptions regarding the presence of correspondences.
@@
Message Passing (MP): Graph neural networks [81, 9, 40] introduced the concept of message passing to DL, where graph nodes are updated depending on received messages from connected nodes. MP was successfully applied, e.g., to various problems in bioinformatics, material science, and chemistry [110, 72].
@@
Continual Learning (CL): The field of CL [93, 73, 92, 61, 8, 102] studies the ability of an ML model to acquire, update, and accumulate knowledge in an incremental manner. This enables models, e.g., to adapt to data-distribution shifts and entirely new tasks. CL is also relevant when frequent re-training is infeasible. Three CL scenarios of increasing difficulty have been proposed [97]: task-incremental (TI), domain-incremental (DI), and class-incremental (CI) CL. The definition of the CI scenario is not limited to classification tasks, but is characterized by the ability of a model to infer which task it is presented with. Prominent approaches for CL include regularization-based methods [41, 46], parameter-isolation methods [54, 80], and replay methods [82, 71, 51, 76]. In contrast to biological neural networks, replay in ML is used for regularizing ML models, where the idea is to learn from new data in tandem with data that is stored in a memory or generated during training. Replay methods can even lead to optimal CL in case of a perfect memory [42].
@@
The notion of knowledge retention and mitigation of forgetting [27] plays a major role in learning. Forgetting is affected by many aspects such as capacity, plasticity, and the quality as well as diversity of training data. The catastrophic forgetting problem in CL [56, 18, 55, 19, 41] refers to learnable parameters adapting to the most recently seen training examples, thereby discarding knowledge from old training examples.
Few-Shot Learning (FSL): Many DL approaches lack the ability to learn efficiently when data is limited. This led to advances in the field of FSL [33], i.e., learning from only a few labeled examples. Available examples are also referred to as support samples. Some methods approach FSL by the means of meta learning [12, 69, 22], where an outer optimization loop learns a strategy for an inner loop, which further adjusts a model to a specific task. The outer loop is performed using large datasets, whereas the inner loop has access to support samples. Another approach to FSL is transfer learning via fine-tuning [15, 60, 100, 1, 45]. Starting from a pre-trained model, learnable parameters or parts thereof are fine-tuned to a specific task using available support samples. Initializing models via pre-training on similar data and tasks typically benefits finding models with high performance after fine-tuning.
Privacy: Common application domains for CL and FSL are autonomous driving and medicine, where data distribution shifts [67] over time are common [87, 64] and data is often scarce. However, storing any kind of personally identifiable data such as patient information in a memory can raise privacy concerns. Even storing hidden representations [29] or detailed, sample-specific labels, raises data-privacy issues due to the possibility of reconstruction attacks [17, 85]. The field of data-free CL [46] has emerged as a viable extension when privacy-preservation is required.
@@ 3 Method
We now describe our parameter-free transduction method in more detail (also see Fig. 1). First, we provide an overview of our method and describe the individual components of our model, which consists of a frozen feature extractor, a memory, and a transduction module (Sec. 3.1). Given extracted features of a query input, we perform transductive reasoning by searching for similar features within memory samples and retrieving stored labels (Sec. 3.2). Finally, we take advantage of intra-query correlations via message passing (Sec. 3.3) and further improve predictions by applying test-time augmentation (Sec. 3.4).
@@ 3.1 General Setup of PARMESAN
Feature Extraction: Our method requires a pre-trained feature extractor $F$. Let $F$ be a neural network with learnable parameters $\theta$ and consisting of an encoder-like or encoder-decoder-like architecture (e.g., U-Net, ConvNeXt-v2, SwinUNETR, etc.) that delivers a feature pyramid $(h)^{l}$ of hidden representations with levels $l=1,...,n$. We refer to $h^{1}$ as the hidden representation of the first level and $h^{n}$ as the hidden representation of the final level (bottleneck). Individual “pixels” in $(h)^{l}$ with their respective features are referred to as nodes. The total number of nodes for level $l$ is $p^{l}=\prod_{dim}res_{dim}^{l}$, where $dim$ refers to the data dimension and $res^{l}$ to the grid resolutions. Our proposed method is designed with the ability to perform tasks like CL and FSL with frozen parameters $\theta$. One can either pre-train $F$ in a supervised or unsupervised manner, or can use existing pre-trained models. We recommend pre-training $F$ on a dense prediction task in order to enrich features with dense information.
@@
Parameter-free Transduction: Let $X$, $H$, and $Z$ be the input-, latent-, and output spaces, respectively. Densely labeled data is referred to as $(x,z)$ with inputs $x\in X$ and labels $z\in Z$. We define a solver $S:X\rightarrow Z$ as the combination of a feature extractor $F:X\rightarrow H$, a memory $M$, and a parameter-free transduction module $G:H\rightarrow Z$. Given a query input $x_{q}$ and a memory $M$ that stores $m$ labeled samples, $G$ then predicts $y_{q}=G(F(x_{q};\theta),M)$. In essence, $G$ performs hierarchical correspondence matching between a query feature pyramid $(h)_{q}^{l}$ and the memory content, followed by label-retrieval and reduction. We use $M$ for both learning and long-term knowledge retention. Knowledge in $M$ can be modified explicitly, allowing for nuanced and sensible management of what to learn, what to retain, and what to forget. Note that this setup shifts the plasticity-robustness trade-off from learnable parameters $\theta$ with parameter-induced forgetting to $M$ with memory-induced forgetting.
@@
Assumptions: We assume grid-based data structures such as images, volumes, videos, and time series. Although we focus on orthogonal grids in this work, our method can be transferred to irregular, adaptive, and isometric grids. We do not impose any restrictions on label-space growth, such as requiring to pre-define the number of output channels, the number of dimensions, or the numerical data type. When taken to extremes, individual tokens of labels can even represent entire data structures such as sets, series, or graphs.
Privacy Preservation: Our method also extends to privacy-preserving settings, where privacy leaks can be mitigated by storing generated [82] or distilled [103, 28] data in $M$. We do not impose strict requirements for having knowledge about labels, e.g., class segments, and their distribution, e.g., class abundances. This setup even allows for keeping memory labels private if needed.
3.2 Hierarchical Correspondence Matching

### v2

@@ Abstract
This work addresses flexibility in deep learning by means of transductive reasoning. For adaptation to new data and tasks, e.g., in continual learning, existing methods typically involve tuning learnable parameters or complete re-training from scratch, rendering such approaches unflexible in practice. We argue that the notion of separating computation from memory by the means of transduction can act as a stepping stone for solving these issues. We therefore propose PARMESAN (parameter-free memory search and transduction), a scalable method which leverages a memory module for solving dense prediction tasks. At inference, hidden representations in memory are being searched to find corresponding patterns. In contrast to other methods that rely on continuous training of learnable parameters, PARMESAN learns via memory consolidation simply by modifying stored contents. Our method is compatible with commonly used architectures and canonically transfers to 1D, 2D, and 3D grid-based data. The capabilities of our approach are demonstrated at the complex task of continual learning. PARMESAN learns by 3-4 orders of magnitude faster than established baselines while being on par in terms of predictive performance, hardware-efficiency, and knowledge retention.
@@
Transduction Memory Correspondence Matching Fast Learning Continual Learning
@@ 1 Introduction
Figure 1: Overview of our parameter-free transduction method for dense prediction tasks. A feature extractor $F$ is applied to a query $x_{q}$ to obtain hidden representations $(h)_{q}^{l}$. A memory $M$ stores labeled samples along with their representations $(h)_{j}^{l}$. Our approach then performs a hierarchical search in $M$, where the objective is to find globally and locally similar nodes w.r.t. $x_{q}$. In each level $l$, we keep the top-$k$ most similar nodes and retrieve their child nodes in level $l-1$. Labels from most similar nodes are retrieved to obtain a raw prediction. Then, message passing (MP) is used to get a final prediction $y_{q}$.
@@
For training task-specific deep learning (DL) models, both data and task are usually well-defined in advance. While this setup holds true in most cases, a major issue emerges when already existing models are required to adapt to new data and new tasks. Satisfying such a flexibility requirement is non-trivial, increasing the difficulty for solving a given problem. Although research in continual learning (CL) resulted in promising approaches, they still suffer from certain restrictions [9, 91] or focus on specific problem-niches [60]. Moreover, proposed solutions are often unstable [7, 40, 39] and require domain experts for deployment and maintenance.
@@
For example, a well established CL approach is replay training, which uses a memory module for knowledge retention [75, 63, 46, 68]. However, replay training turns out to be challenging in practice due to several reasons. First, models require high plasticity to learn and unlearn continually, which in turn limits their ability to retain knowledge. Forcing models to be robust, i.e., retain knowledge via replay training, is thus somewhat contradicting the plasticity requirement for learning new concepts. Second, all relevant knowledge has to be incorporated into model parameters via training, thereby rendering the available memory capacity irrelevant for inference. Third, the balancing of memory samples and new samples for training is non-trivial in practice and often leads to memory overfitting [46, 6, 90]. Finally, replay does not allow to unlearn specific examples easily, e.g., in case they are not needed anymore. We argue that simplicity and flexibility are not necessarily mutually exclusive properties of modern DL approaches. In contrast to common, induction-based approaches, where predictions are inferred from general, learned principles, we embrace transduction, which is characterized by reasoning from specific training cases to specific test cases [19].
@@
Our Work: We propose PARMESAN (parameter-free memory search and transduction), an approach that separates computation from memory and extends transduction to dense prediction tasks while at the same time retaining high flexibility. Our method combines several concepts to perform hierarchical correspondence matching with a memory that contains task-specific, labeled data. Rather than requiring to incorporate all knowledge into learnable parameters, the memory in our method extends the capacity of existing models. We further propose a message passing approach to take advantage of correlations within a query sample, as well as a sparsity approach to retain memory efficiency.
@@
PARMESAN is parameter-free in the sense that it does not have learnable parameters. Instead, learning can be easily performed via memory consolidation, i.e., adding, removing, or modifying memory samples. Our method allows for fast learning and unlearning of individual examples or parts thereof. Unlike most CL methods, our approach does not require any continual parameter-training by design. It therefore neither suffers from parameter-induced forgetting and memory overfitting, nor does it require complicated and energy-hungry training.
@@
PARMESAN can handle various data dimensionalities such as 1D sequences, 2D images, 3D volumes, as well as spatio-temporal data. We do not impose major restrictions on memory size and allow flexible memory management. Our method can be applied in combination with common DL architectures, thereby enhancing learning speed. This is relevant, e.g., in test-time learning or when dealing with limited computational resources. It is also task-agnostic, i.e., it transfers to different dense prediction tasks as well as multi-tasking.
@@
  • A parameter-free transduction method for dense prediction tasks using hierarchical correspondence matching with a memory.
  • A parameter-free message passing approach to take advantage of intra-query correlations.
  • A novelty-sparsity approach to achieve high memory diversity while remaining hardware-efficient.
  • A demonstration of our method’s flexibility by using it in combination with common DL archictectures for semantic segmentation and depth estimation using the Cityscapes and JSRT Chest X-ray datasets.
  • We show that our method can be successfully applied to CL, being on par with well-established baselines in terms of predictive performance and memory efficiency while at the same time learning by 3-4 orders of magnitude faster and having stable knowledge retention.
@@ 2 Related work
Our method is a novel combination of concepts from the following fields:
@@
Memory Networks: Recently, memory networks experience renewed interest [27, 94, 22, 23, 62, 55, 1]. The notion of separating computation and memory [23] allows incorporating new information easily while at the same time enabling treatment of memory contents as variables. Memory consolidation [20, 3] concerns the strategy for deciding which knowledge to keep in the presence of limited memory resources. A sophisticated memory consolidation method is novelty search [38], where only the most informative examples are kept. We expand on this approach when introducing sparsity to stored samples.
@@
Transduction: Recently, transduction methods experience a revival due to their effectiveness in computer vision [2]. A well-established transduction method is k-nearest neighbors ($k$-NN) [14, 32], which aims to find similar examples w.r.t. common features. Transduction via $k$-NN was demonstrated to work well for image classification [54], where labeled training samples and their hidden representations are stored in memory. Moreover, $k$-NN in combination with vision transformers [89, 11] yielded promising results for dense out-of-distribution detection [18]. However, making $k$-NN scalable to dense prediction tasks at high resolutions remains underexplored.
@@
Correspondence Matching: Dense correspondence matching [67, 93, 43, 48, 31, 28, 78, 12] is a fundamental problem in computer vision, often acting as a starting point for downstream tasks like optical flow and depth-estimation. In this context, promising methods to improve feature correlation have been proposed [85, 29, 50]. Also, feature pyramids were exploited for object detection [21, 42] and semantic feature matching [86, 98]. Recently, foundation models [4] have been applied for feature matching in one-shot semantic segmentation [44]. Many existing matching approaches use a single reference sample with known correspondences, but typically can not handle more than one reference sample. Moreover, most methods rely on specific assumptions commonly found in setups with videos or stereo / multi-view images. In contrast, our proposed method allows matching of arbitrary samples and does not require strong assumptions regarding the presence of correspondences.
@@
Message Passing (MP): Graph neural networks [74, 10, 35] introduced the concept of MP to DL, where graph nodes are updated depending on received messages from connected nodes. MP was successfully applied, e.g., to various problems in bioinformatics, material science, and chemistry [97, 64].
@@
Continual Learning (CL): The field of CL [84, 65, 83, 56, 9, 91, 96] studies the ability of an ML model to acquire, update, and accumulate knowledge in an incremental manner. This enables adaptation to data-distribution shifts [61, 79] and new tasks. CL is also relevant when frequent re-training is infeasible. Three CL scenarios of increasing difficulty have been proposed [88]: task-incremental (TI), domain-incremental (DI), and class-incremental (CI). The definition of the CI scenario is not limited to classification, but is characterized by the ability of a model to infer which task it is presented with.
Prominent CL approaches are based on regularization [36, 41], parameter-isolation [49, 72], and replay [75, 63, 46, 68, 37]. In contrast to biological neural networks, replay in ML is used for regularization, where the idea is to learn from new data in tandem with data that is stored in a memory.
@@
In CL, the notion of knowledge retention [24] plays a major role in learning. Forgetting is affected by aspects such as capacity, plasticity, and the quality as well as diversity of training data. The catastrophic forgetting problem in CL [52, 16, 51, 17, 36] refers to learnable parameters adapting to the most recently seen training data, thereby discarding knowledge from old data. Note that storing sensitive data such as patient information in a memory raises privacy concerns due to the possibility of reconstruction attacks [15, 77]. The field of data-free CL [41] using generated [75] or distilled [92, 25] data has emerged as a viable extension when privacy-preservation is needed.
@@ 3 Method
In this section we describe our parameter-free transduction method (cf. Fig. 1) in more detail. First, we provide an overview of our method and describe the individual components of our model, which consists of a frozen feature extractor, a memory, and a transduction module (Section 3.1). Given a query input, we perform dense hierarchical correspondence matching with memory contents and take advantage of intra-query correlations via message passing (Section 3.2). Finally, we improve memory efficiency by introducing sparsity to stored samples (Section 3.3).
@@ 3.1 General Setup of PARMESAN
Our proposed method requires a pre-trained feature extractor $F$. Let $F$ be a neural network with learnable parameters $\theta$ consisting of an encoder-like or encoder-decoder-like architecture (e.g., [69, 73, 80]) that delivers a feature pyramid $(h)^{l}$ of hidden representations with levels $l=1,...,n$. We refer to $h^{1}$ as the hidden representation of $l=1$ (high-res) and $h^{n}$ as the hidden representation of $l=n$ (low-res). Individual “pixels” in $(h)^{l}$ with their respective features are called “nodes”. The total number of nodes for level $l$ is $p^{l}=\prod_{dim}res_{dim}^{l}$, where $dim$ refers to the data dimension and $res^{l}$ to the grid resolutions. Our method is designed to perform CL with frozen parameters $\theta$. $F$ can be pre-trained in a supervised or unsupervised manner, or can be taken from public repositories. Although not strictly required, we recommend pre-training $F$ on a dense prediction task in order to enrich features with dense information.
@@
Let $X$, $H$, and $Z$ be the input-, latent-, and output spaces, respectively. Densely labeled data is referred to as $(x,z)$ with inputs $x\in X$ and labels $z\in Z$. We define a solver $S:X\rightarrow Z$ as the combination of a feature extractor $F:X\rightarrow H$, a memory $M$, and a parameter-free transduction module $G:H\rightarrow Z$. Given a query input $x_{q}$ and a memory $M$ that stores $m$ labeled samples, $G$ predicts $y_{q}=G(F(x_{q};\theta),M)$. $G$ performs hierarchical correspondence matching between a query feature pyramid $(h)_{q}^{l}$ and memory contents, followed by label-retrieval and postprocessing. We use $M$ for both learning and long-term knowledge retention. Knowledge in $M$ can be modified explicitly, allowing for nuanced and sensible management of what to learn, what to retain, and what to forget. Note that this setup shifts the plasticity-robustness trade-off from learnable parameters with parameter-induced forgetting to a memory with memory-induced forgetting.
In contrast to other approaches, our method is designed to learn fast in the sense that it does not require adapting $\theta$ to perform complex tasks such as CL. Our method only requires memory consolidation, which in our case means saving examples in $M$ and extracting their feature pyramids $(h)_{j}^{l}$ in a single forward pass with frozen $\theta$.
@@
We assume grid-based data structures such as images, volumes, and time series. Although we focus on orthogonal grids in this work, our method can be transferred to irregular, adaptive, and isometric grids. We do not impose any restrictions on label-space growth, such as requiring to pre-define the number of output channels, the number of dimensions, or the numerical data type. When taken to extremes, individual tokens of labels can even represent entire data structures such as sets, series, or graphs.
3.2 Correspondence Matching and Intra-Query Transduction